solve the cubic equation 2z3 -5z2+z-5=0


1
Expert's answer
2022-06-07T00:22:31-0400
"a z^3 +b z^2 +c z+d=0\\
2 z^3 - 5 z^2 + z - 5=0"

"a=2,\; b=-5,\; c=1,\; d=-5"

"\Delta_0=b^2-3ac=19\\
\Delta_1=2b^3-9abc+27a^2d=-700"

"C=\sqrt[3]{\frac{\Delta_1+\sqrt{\Delta_1^2-4\Delta_0^3}}{2}}=\sqrt[3]{-350+3\sqrt{12849}}"

"z_k=-\frac{1}{3a}(b+\xi^kC+\frac{\Delta_0}{\xi^kC}),\; k=0,1,2"

"\xi=\frac{-1+i\sqrt{3}}{2}"

"z_0\!=\!\frac{5}{6}\!+\!\frac{1}{6} ( (350\! -\! 3 \sqrt{12849})^{1/3} 
\!+\! (350 \!+ \!3 \sqrt{12849})^{1/3})"

"z_1=\frac{5}{6} -\frac{1}{12} (1 + i \sqrt{3}) (350 - 3 \sqrt{12849})^{1/3} \\
- \frac{1}{12} (1 - i \sqrt{3}) (350 + 3 \sqrt{12849})^{1/3}"

"z_2=\frac{5}{6} -\frac{1}{12} (1 - i \sqrt{3}) (350 - 3 \sqrt{12849})^{1/3} \\
- \frac{1}{12} (1+ i \sqrt{3}) (350 + 3 \sqrt{12849})^{1/3}"